OpenStudy (anonymous):

Which improper fraction is equivalent to 2 3/9? 6n+4 4-6n 4n-6 6+4n

hartnn (hartnn):

2 3/9 = 2+3/9 = (18+3)/9 = 21/9 or 7/3 or 2 1/3